# ZeroMerge: Parameter-Free KV Cache Compression for Memory-Efficient Long-Context LLMs

### v1

@@ 3.4 Residual Token Merging
The residual component dynamically consolidates evicted tokens through similarity-driven aggregation. When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we:
@@
Algorithm 1 ZeroMerge Online Compression
1:  Input: Budgets $B_{c}$, $B_{r}$, $B_{p}$, decay $\lambda$
@@
12:           $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
13:           if $|\mathbf{K}_{r}|>B_{r}$ then
14:              Merge via (6)–(7)
15:           end if
@@ 3.5 Attention Output Stabilization
$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{s=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{s}/\sqrt{d}+\alpha\log w_{s}\right)},$ (8)
@@
where $w_{j}$ denotes the fusion count of token $j$ (with $w_{j}=1$ for uncompressed tokens), and $\alpha\in(0,1]$ controls residual compensation intensity. This formulation addresses two key challenges:
  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors through Eq. 7, creating a mismatch between their key vectors ($\mathbf{k}_{r}$) and the original value distribution. The $\log w_{j}$ term compensates for this representation shift.
  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, preventing suppression of critical context.
@@
This theoretical guarantee ensures uncompressed tokens maintain their relative attention dominance despite cache compression. The compensation mechanism effectively:
  • Preserves attention mass for critical context tokens
  • Prevents over-amplification of merged token scores
4 Experiment
In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction. Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used. Secondly, we verify the effects of our method on memory control and throughput. Moreover, we compare the generation quality of different methods in the field of long texts and tasks. Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.
@@ 4.1 Experiment setup
The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA-2 7B, LLaMa3-8B-Instruct, and Mistral-7B-Instruct. The experimental equipment is an NVIDIA A800-80GB GPU. For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs. It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
@@
  • Full KV (FKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
@@
  • ZeroMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based 0-shot merging. Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens. Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
4.2 Inference efficiency and resource consumption
Figure 3: VRAM Utilization and Decoding Throughput Across Sequence Lengths. ZeroMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating OOM via dynamic KV cache compression.
We evaluate the long - context processing efficiency on a single NVIDIA A800 GPU with 80GB VRAM (Video Random Access Memory). The experiment uses the LLaMA2 - 7B model, which has 32 layers, 32 attention heads, and an embedding dimension of 4096, in float32 precision. The benchmark is composed of a synthetic QA task, where the cache tokens range from 20K to 54K. For the baseline measurements, we use full KV caching without compression. In contrast, the ZeroMerge method enforces a token cache budget of 18K.
The original Llama2-7B model (float32) required 25GB VRAM for parameter loading alone. During prefill phase, processing 20K prompt tokens generated 20GB KV cache across 32 layers, totally consuming 56% of total available VRAM (Fig.3). In baseline inference, KV cache grew linearly with sequence length at 1MB/token. By 54K tokens, cache memory dominated the mast of VRAM, triggering OOM errors. Concurrently, decoding throughput degraded from 9 tokens/sec (20K context) to 4 tokens/sec (52K context)—due to the increased computation of attention score.
By enforcing a 18K token cache ceiling through two-phase compression (attention sink preservation + similarity-based merging), ZeroMerge: Reduced prefill-phase KV cache to 18GB (10% reduction); Maintained constant 43GB VRAM usage during decoding (vs. baseline’s 45 to 79 GB linear growth); Sustained stable throughput at 9 tokens/sec across 20K–54K token contexts; Eliminated OOM constraints through dynamic cache compression.
4.3 Workload-Scalable KV Efficiency
We evaluate ZeroMerge against three baseline methods (FullKV, H2O, LESS) on an NVIDIA A800 GPU (80GB VRAM) using LLaMA2-7B (float32) with diverse workloads: sequence lengths (2K–54K tokens), batch sizes (4–16), and document-QA tasks. Each method is tested under identical decoding constraints (temperature=0.0, greedy sampling) to isolate computational overhead. Metrics include throughput (tokens/sec), latency (sec), and OOM thresholds.
Table 1: Workload-Scalable KV Cache Compression: ZeroMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
@@ 4.3 Workload-Scalable KV Efficiency
FullKV H2O (5%) Less (5%) ZeroMerge (5%)
@@
As demonstrated in Table 1, ZeroMerge consistently outperforms existing methods across diverse workload configurations. At moderate batch sizes (batch=4, 4K+4K tokens), ZeroMerge achieves 23.4% higher throughput (77.0 vs 62.4 tokens/sec) and 18.9% lower latency (212.7s vs 262.1s) compared to FullKV, while H2O and LESS suffer 29.7% and 75.9% performance degradation respectively. Under high-throughput demands (batch=8, 4K+4K tokens), ZeroMerge successfully processes requests where both H2O and LESS encounter OOM errors, delivering 124% throughput improvement (146.7 vs 65.5 tokens/sec) over FullKV. Most notably, at large batch sizes (batch=16, 2K+2K tokens), ZeroMerge sustains 2.1× higher throughput (281.9 vs 133.1 tokens/sec) with 52.8% latency reduction (116.2s vs 246.2s), showcasing its unique capability to decouple computational load from sequence length. These results validate ZeroMerge’s dual advantage, eliminating memory bottlenecks through strict cache budgeting.
4.4 Inference quality
@@
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. The study compares three KV cache compression methods across two 7B models (Llama2, Falcon) at 20%, 10%, and 5% cache budgets. H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network. ZeroMerge operates with $\alpha=0.6$ (influence ratio of residual cache ) and $\lambda=0.98$ (the decay rate of the historical contribution) without task-specific training.
@@
Our evaluation demonstrates that ZeroMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. When compressing the KV cache to 5% of its original size, ZeroMerge maintains near-original performance on Llama2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZeroMerge’s performance preservation significantly outperforms both baselines.
@@
On Falcon-7B, a model with distinct architectural characteristics, ZeroMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation—likely due to its reliance on C4 dataset training—ZeroMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
@@
The results highlight two critical trends: First, higher compression ratios amplify ZeroMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
@@
method Llama2 7B Falcon 7B
FULL 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
@@
5% ZeroMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
Acknowledgements
Do not include acknowledgements in the initial version of the paper submitted for blind review.
If a paper is accepted, the final camera-ready version can (and usually should) include acknowledgements. Such acknowledgements should be placed at the end of the section, in an unnumbered section that does not count towards the paper page limit. Typically, this will include thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ Acknowledgements
Authors are required to include a statement of the potential broader impact of their work, including its ethical aspects and future societal consequences. This statement should be in an unnumbered section at the end of the paper (co-located with Acknowledgements – the two may appear in either order, but both must be before References), and does not count toward the paper page limit. In many cases, where the ethical impacts and expected societal implications are those that are well established when advancing the field of Machine Learning, substantial discussion is not required, and a simple statement such as the following will suffice:
“This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.”
The above statement can be used verbatim in such cases, but we encourage authors to think about whether there is content which does warrant further discussion, as this statement will be apparent if the paper is later flagged for ethics review.

### v2

@@ 3.4 Residual Token Merging
The residual component dynamically consolidates evicted tokens through similarity-driven aggregation. When merging a candidate token $(\mathbf{k}_{t},\mathbf{v}_{t})$ into $\mathbf{K}_{r}$, we adopt the following three-step procedure:
@@
Algorithm 1 ZSMerge Online Compression
1:  Input: Budgets $B_{p}$, $B_{c}$, $B_{r}$, decay $\lambda$
@@
12:           if $|\mathbf{K}_{r}|+1\leq B_{r}$ then
13:              $\mathbf{K}_{r}\leftarrow\mathbf{K}_{r}\cup\{\mathbf{k}_{\hat{c}}\}$, $\mathbf{V}_{r}\leftarrow\mathbf{V}_{r}\cup\{\mathbf{v}_{\hat{c}}\}$
14:           else
15:              Merge token $(\mathbf{k}_{\hat{c}},\mathbf{v}_{\hat{c}})$ into $\mathbf{K}_{r}$ via (6)-(7)
@@ 3.5 Attention Output Stabilization
$\hat{a}_{t}^{(T)}$ $=\frac{\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{t}/\sqrt{d}+\alpha\log w_{t}\right)}{\sum_{i=1}^{T}\exp\left(\mathbf{q}_{T}^{\top}\mathbf{k}_{i}/\sqrt{d}+\alpha\log w_{i}\right)},$ (8)
@@
where $w_{i}$ represents the fusion count of token $i$ (with $w_{i}=1$ for uncompressed tokens). The hyperparameter $\alpha\in(0,1]$ regulates residual compensation intensity, with empirical analysis suggesting an optimal range of $\alpha\in[0.6,0.8]$ for balanced performance. Notably, setting $\alpha=0$ reduces the merging algorithm to a degenerate case equivalent to a pure eviction policy, eliminating residual compensation entirely.
With Eq. 8, we address two key challenges:
  • Representation Bias Correction: Merged tokens aggregate multiple historical vectors via Eq. 7, creating a mismatch between their key vectors ($\mathbf{k}_{r}$) and the original value distribution. The $\log w_{j}$ term compensates for this representation shift.
  • Attention Mass Conservation: The compensation term preserves the relative attention mass between compressed and uncompressed tokens, ensuring residual compensation does not suppress critical uncompressed components.
@@
This theoretical guarantee ensures that uncompressed tokens retain their relative dominance in attention allocation despite cache compression, even as the denominator accounts for upper-bounded contributions from compressed tokens. The compensation mechanism effectively preserves attention mass for critical context tokens while preventing over-amplification of merged token scores, thereby maintaining the integrity of the original attention distribution under compression.
4 Experimental Evaluations
In this chapter, we aim to verify the effectiveness of our proposed method in terms of performance preservation and memory reduction. Firstly, we introduce the experimental setup, including the selected model architectures, the baseline methods for comparison, and parameters such as the datasets used. Secondly, we verify the performances of our method on memory control and throughput. Moreover, we compare the generation quality of different methods in the field of long texts and tasks. Finally, we discuss the trade-off between memory, time, and performance as well as the impact of relevant hyperparameters on the model.
@@ 4.1 Experiment Setup
The backbone models selected in the experiment are some small-scale open-sourced LLMs, including LLaMA2-7B, LLaMA3-8B-Instruct, and Mistral-7B-Instruct. The experimental equipment is an NVIDIA A800-80GB GPU. For the dataset, we choose LongBench, which is a comprehensive benchmark dataset specifically designed for evaluating the performance of different methods on tasks that involve long-context inputs. It contains a wide variety of natural language processing tasks with different lengths of text sequences, covering multiple domains such as news, novels, and academic papers, thus providing a rich and diverse set of test cases for our experiments.
@@
  • Full KV (FullKV): Caches all keys and values per token in each layer, serving as the vanilla baseline with maximal computational load but theoretically optimal performance.
@@
  • ZSMerge (Ours): Builds upon merging paradigm but eliminates parametric dependencies through heuristic-based zero-shot merging. Strictly enforces preset cache bounds via two-phase compression: 1) Attention sink preservation for initial/recent tokens, 2) Distance-aware similarity clustering for intermediate tokens. Minimizes embedding distortion through revised attention scoring that accounts for merged token positional densities, achieving better generalization across compression ratios than prior learning-based methods.
4.2 Inference Efficiency Gain
To demonstrate the benefits of ZSMerge in improving inference efficiency, we conducted two proof-of-concept experiments. The first (Figure 3) compares the performance of full KV caching and ZSMerge under increasing sequence lengths in a specific inference case. The second (Table 1) evaluates ZSMerge against full KV caching and other baseline methods under various workloads, including different sequence lengths, batch sizes, and model sizes. Below, we summarize the memory and throughput improvements achieved by ZSMerge.
Figure 3: VRAM Utilization and Decoding Throughput Across Sequence Lengths. ZSMerge enforces constant memory footprint (43GB) and sustains 9 tokens/sec decoding rate beyond 54K tokens, eliminating out-of-memory (OOM) via dynamic KV cache compression.
Table 1: Workload-Scalable KV Cache Compression: ZSMerge Outperforms Baselines in Throughput (tokens/sec) and Latency (seconds) Across Sequence Lengths and Batch Sizes.
@@ 4.2 Inference Efficiency Gain
FullKV H2O (5%) Less (5%) ZSMerge (5%)
@@
4.2.1 Memory Reduction
Specific Case Analysis
In the baseline setup, the LLaMA2-7B model required 25GB of VRAM for parameter loading, with an additional 20GB KV cache generated during the prefill phase for 20K tokens. This linear growth in KV cache size, at 1MB per token, led to OOM errors as sequence lengths approached 54K tokens.
ZSMerge, constrained by an 18K token cache budget, reduced KV cache size by 10% during the prefill phase and maintained VRAM usage at a constant 43GB during decoding. This prevented the baseline’s linear memory growth (up to 79GB) and completely eliminated OOM errors, enabling efficient long-context processing.
4.2.2 Throughput Improvement
Micro-Level Analysis
Decoding throughput for the baseline dropped from 9 tokens/sec at 20K tokens to 4 tokens/sec at 54K tokens due to increasing attention computation overhead. ZSMerge, in contrast, maintained a consistent throughput of 9 tokens/sec across the same sequence length range by dynamically merging less relevant tokens while preserving critical attention information.
Macro-Level Analysis
Under diverse workloads, ZSMerge consistently outperformed the baseline and other methods. For moderate batch sizes (batch=4), it achieved 23.4% higher throughput and 18.9% lower latency than full KV caching. At higher batch sizes (batch=8), ZSMerge handled requests that caused OOM errors for other methods, delivering a 124% throughput improvement. For large batch sizes (batch=16), it sustained 2.1× higher throughput and halved latency compared to full KV caching.
In summary, ZSMerge effectively reduces memory usage and maintains high throughput across various workloads, outperforming full KV caching and other baselines. Its dynamic cache compression mechanism ensures scalability and efficient long-context processing without sacrificing performance.
4.3 Inference Quality Loss
@@
We evaluate text generation quality on the XSum abstractive summarization dataset (16k news articles with single-sentence summaries) using ROUGE-1/2/L metrics. The experiment compares three KV cache compression methods across two 7B models (LLaMA2-7B and Falcon) at 20%, 10%, and 5% cache budgets. H2O balances recent tokens and heavy hitters per its original design, while LESS employs its recommended C4-trained merging network. For ZSMerge, the hyperparameters were set to static values based on empirical validation: the residual compensation intensity ratio ($\alpha$) was fixed at $0.6$ to balance performance across tasks, and the historical contribution decay rate ($\lambda$) was set to $0.98$ to ensure stable context propagation while prioritizing recent attention patterns. These values were selected through comparative analysis against other configurations (e.g., $\alpha\in[0.6,0.8]$ and $\lambda=0.98$ showed optimal trade-offs between compression efficiency and generation quality) and required no task-specific training.
@@
Our evaluation demonstrates that ZSMerge achieves superior text generation quality across varying compression ratios and model architectures compared to existing KV cache compression methods. When compressing the KV cache to 5% of its original size, ZSMerge maintains near-original performance on LLaMA2-7B, even slightly exceeding the uncompressed baseline in ROUGE scores, while H2O and LESS exhibit measurable degradation under the same conditions. This advantage becomes more pronounced at higher compression ratios, where ZSMerge’s performance preservation significantly outperforms both baselines.
@@
On Falcon-7B, a model with distinct architectural characteristics, ZSMerge demonstrates stronger generalization than parameter-dependent approaches. While LESS suffers severe quality degradation-likely due to its reliance on C4 dataset training-ZSMerge retains over half of the baseline performance even at extreme 5% compression. In contrast, H2O struggles to balance recent and heavy-hitter tokens effectively, particularly for Falcon’s attention patterns.
@@
The results highlight two critical trends: First, higher compression ratios amplify ZSMerge’s relative advantages, as its heuristic merging mechanism better preserves semantically critical tokens compared to purely eviction policies. Second, architecture-agnostic design enables consistent cross-model robustness, avoiding the domain adaptation challenges faced by data-driven methods like LESS.
@@
method LLaMA2-7B Falcon-7B
Full KV 30.59 / 11.34 / 25.50 27.06 / 8.79 / 22.39
@@
5% ZSMerge 30.60 / 11.67 / 25.72 15.04 / 3.29 / 12.73
5 Conclusion
This study addresses the critical challenge of memory and computational inefficiency in long-context language models through ZSMerge, a dynamic KV cache compression framework. By introducing head-level memory allocation, residual merging with compensated attention scoring, and architecture-agnostic adaptation, the method achieves sublinear memory growth while preserving generation quality across diverse workloads. Empirical validation demonstrates that ZSMerge reduces VRAM consumption by 82% at 54K tokens,triples inference efficiency under a 5% compression ratio, and outperforms eviction-based baselines by 12-34% in text quality metrics—all without task-specific training. The framework advances practical deployment of long-context LLMs by fundamentally decoupling memory constraints from sequence length, enabling efficient processing of extended dialogues and documents while maintaining semantic coherence.
@@ 5 Conclusion
This paper presents work whose goal is to advance the fields of machine learning and sustainable artificial intelligence. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.